# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 7.3 Proofs of Section˜5
$=\frac{1}{n}\sum_{i=1}^{n}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\notin A}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\notin A}.$
@@
$\leq(1+a)\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]+\frac{(1+3/a)D\log(n)^{6}}{3n}\log\Big{(}\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{2}}\Big{)}$
$\quad+(1+a)\tau.$
Like before, we separate $a\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$. Using Lemma˜2.1 again, we obtain
$\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}$
$\quad-\int_{\mathbb{R}^{d}\setminus A}\int\int|\tilde{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y.$
@@
Collecting all other terms, inserting $\log(\sigma_{\min})\sim\log(n)$, setting $\delta_{1}=\delta_{2}=\frac{1}{3n}$ and using the union bound leads to the result. The infimum can be used due to the presence of other nonzero terms. ∎
Preparation for the proof of Theorem˜5.3.
The proof of Theorem˜5.3 requires some additional results to bound the covering number and apply the approximation result. We further need to look carefully into the approximation result by Gühring et al. (2020).
@@
Hence we can bound
$\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\leq\mathcal{N}(\frac{\tau}{\log(n)^{4}},\mathcal{M},\|\cdot\|_{\infty}).$
From Suzuki (2019, Lemma 3) we know that the covering number of the set ReLU networks of a $0,1]^{d}$, denoted by $\mathsf{NN}_{\text{ReLU}}$ cube is
@@
To consider functions on $\log(n),\log(n)]^{d}\times[0,1]$, we proceed analogously to (Yakovlev and Puchkin, 2025, p.46), multiplying the first $d$ coordinates of the weight matrix of the first lemma with $\log(n)$ and dividing the input vector with the same value. This has of course an impact on the bound of the weights, which will scale up by the factor $\log(n)$. Additionally, the weights of the clipping layers are of order $\log(n)^{3}$. We thus obtain
@@
The network $f_{\mathsf{NN}}$ can be chosen with at most
$L\leq c\cdot\log\left(\varepsilon^{-\frac{\xi}{\beta-1}}\right)$
hidden layers and
$S\leq c\cdot\varepsilon^{-\frac{d}{s-1}}\cdot\log^{2}\left(\varepsilon^{-\frac{s}{s-1}}\right)$
nonzero weights, where $c$ is a constant.
@@
As $v$ is in $C^{\infty}$, we can apply (Gühring et al., 2020, Corollary 4.2) for arbitrary large $s$. However, the use of a larger $s$ will lead to a larger bound on the absolute value of the partial derivatives up to the order $s$. Since this scales up the approximation error, we need a precise quantification. The next result uses the $\log$-Sobolev inequality to bound higher derivatives of $v$.
@@
The bound on mixed derivatives follows exactly the same lines as the proofs of Lemma˜7.7 and Lemma˜7.8
Now we are ready to proof Theorem˜5.3. Proof of Theorem˜5.3.
We scale the function down by $\sigma_{\min}^{-s-2}\log^{2}(n)$ and approximate $v\cdot\sigma_{\min}^{-s-2}\log^{2}(n)$ instead of $v$. After that, we add another layer to scale the output of the neural net up. The approximation error $\varepsilon$ will also scale up by $\sigma_{\min}^{-s-2}\log^{2}(n)$. The maximal weight in the inner network will then no longer depend on this bound, but the largest weight might increase. Hence we can set
$W=C(\varepsilon^{-\frac{1}{s-2}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)).$
Inserting everything into Theorem˜5.2 and setting $\tilde{d}=d+1$, we obtain for $a<1$ with a probability of $1-\frac{1}{n}$ for $n$ big enough
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}$
$\lesssim+\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t+(2+a)\tau+a\log(n)^{4}+\varepsilon\sigma_{\min}^{-s-2}\log^{2}(n)$
$\quad+\frac{a^{-1}\log(n)^{7}}{n}\varepsilon^{-\frac{\tilde{d}}{s-1}}\log\big{(}2\big{(}\tau^{-1}\log(n)^{7}L(B\vee 1)(\varepsilon^{-\frac{1}{s-1}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)\big{)}\big{)}$
$\quad+(2+a)\tau+a\log(n)^{6}+\varepsilon\sigma_{\min}^{-s-2}\log^{2}(n).$ (36)
Note that $B$ and $L$ will depend on $\varepsilon$, thus we are restricting feasible choices of $\varepsilon$ to choices that grow at most polynomial in $n^{-1}$. Ignoring logarithmic terms, we first solve for $a$:
$\frac{a^{-1}\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\sim a\quad\Longleftrightarrow\quad a=\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}.$
$\tau$ can be set such that the corresponding term is of the same order, it suffices to set
$\tau\sim\frac{\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}}{2+\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}}.$
This choice, as well as the choice of $\varepsilon$, influence the first term in (36) only logarithmically. Now we solve for $\varepsilon$
$\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}\sim\varepsilon\sigma_{\min}^{-s-2}\quad\Longleftrightarrow\quad\varepsilon\sim n^{-\frac{s-1}{\tilde{d}+2s-2}}\sigma_{\min}^{(s+2)\frac{2s-2}{\tilde{d}+2s-2}}.$
Restricting $\sigma_{\min}$ to polynomials in $n$, we ensure $\varepsilon$ is polynomial in $n$. Thus, ignoring logarithmic factors in $n$, we obtain with a probability of $1-\frac{1}{n}$ for $n$ big enough
@@
To choose $\sigma_{\min}$ such that is is optimal in the setting of Theorem˜3.1, we have to set it such that
$\sigma_{\min}^{1+\alpha}\sim\Big{(}n^{-\frac{s-1}{d+2s-2}}\sigma_{\min}^{(s+2)\big{(}\frac{2s-2}{\tilde{d}+2s-2}-1\big{)}}\Big{)}^{\frac{1}{2}}\quad\Longleftrightarrow\quad\sigma_{\min}\sim n^{-\frac{1}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}.$
This choice is a polynomial of $n^{-1}$ and hence a feasible choice for $\sigma_{\min}$. For $n\geq 1$ and $\sigma_{\min}<1$, we know that the network approximates the Lipschitz constant of $v$ with precision $\varepsilon\sigma_{\min}^{-s-2}$. Plugging in the choice of $\sigma_{\min}$ leads to
$\varepsilon\sigma_{\min}^{-s-2}\sim n^{-\frac{2+2\alpha}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}\leq 1.$
Hence we obtain for the Lipschitz constant $\hat{L}_{t}$ of $\hat{v}_{t}$:
$e^{\int_{0}^{1}\hat{L}_{t}\;\mathrm{d}t}\leq e^{\int_{0}^{1}\Gamma_{t}+1\;\mathrm{d}t}\leq e^{C+1}.$
Note that we cannot let $s\rightarrow\infty$, as this will blow up the constant. However, for every fixed $\eta>0$, we can choose $s$ such that
$s=\big{\lceil}(5+2\alpha)\frac{\tilde{d}}{\eta}+1\big{\rceil}.$
This leads to the final choice
$\sigma_{\min}\sim n^{-\frac{1}{\tilde{d}+4\alpha+4+\eta}}$
Now we use the results of Lemma˜7.4, set $\delta=\frac{1}{2n}$ and use the union bound. Combining this with Theorem˜3.1, Theorem˜4.5, (Kunkel and Trabs, 2025b, Theorem 3.3) and $\tilde{d}=d+1$, we obtain with a probability of $1-\frac{1}{n}$
@@
7.4 Additional proofs of Section˜7 and helper results
Proof of Lemma˜7.1.
@@ 7.4 Additional proofs of Section˜7 and helper results
$=\frac{(\int y_{j}(-\frac{x_{i}-t^{\gamma}y_{i}}{\sigma_{t}^{2}})p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)-(\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int(-\frac{x_{i}-t^{\gamma}y_{i}}{\sigma_{t}^{2}})p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)}{(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)^{2}}$
@@
$\frac{\partial}{\partial x_{i}}v^{j}_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ji}.$
Proof of Lemma˜7.2.
@@
$\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}$ $-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}\;\mathrm{d}t$
$\leq\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t+\int_{0}^{t^{*}}\big{|}\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}\;\mathrm{d}t.$
@@
For the second integral, we can bound the Covariance term by
$\big{|}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}$ $\leq C.$
Now we get
@@
$\leq C\int_{0}^{t^{*}}\Big{|}\gamma t^{\gamma-1}\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big{|}\;\mathrm{d}t+C\int_{0}^{t^{*}}\Big{|}\frac{\sigma_{t}^{\prime}t^{2\gamma}}{\sigma^{3}_{t}}\Big{|}\;\mathrm{d}t\leq\vartheta^{2}C\int_{0}^{t^{*}}\gamma t^{\gamma-1}\;\mathrm{d}t+\vartheta^{2}C\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta)).$
Proof of Lemma˜7.4.
@@
$\int_{\mathbb{R}^{d}\setminus A}p^{*}(y)\;\mathrm{d}y$ $\leq\int_{\mathbb{R}^{d}\setminus A}|y|^{2}p^{*}(y)\;\mathrm{d}y=\frac{\int_{\mathbb{R}^{d}\setminus A}|y|^{2}\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}$
$\leq e^{2L}\frac{\int_{\mathbb{R}^{d}\setminus A}|y|^{2}\exp(-\frac{|y|^{2}}{2})\;\mathrm{d}y}{\int\exp(-\frac{|y|^{2}}{2})\;\mathrm{d}y}=e^{2L}\Big{(}2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big{)}^{d}.$
@@
$0\leq\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y=[-y_{1}\exp(-\frac{y_{1}^{2}}{2})]_{\log(n)}^{\infty}-\int_{\log(n)}^{\infty}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}\leq\log(n)\exp(-\frac{\log(n)^{2}}{2}),$
@@
$e^{2L}\Big{(}2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big{)}^{d}$ $\lesssim\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}.$
@@
$\mathbb{E}_{X_{i}}[|B|]\lesssim\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}.$
Now Markov’s inequality yields
$\mathbb{P}(B\geq n^{-\frac{1}{2}})$ $\lesssim\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}},$
$\mathbb{P}(B\leq-n^{-\frac{1}{2}})$ $\lesssim\log(n)^{6}\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}}.$
@@
$\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}}\lesssim\frac{1}{3n}\quad\Longleftarrow\quad n\gtrsim e^{3(d+7+\frac{1}{2})},$
@@
Proof of Lemma˜7.5.
From the proof of Theorem˜5.2 and as $\log(\sigma_{\min}^{-1})\sim\log(n)$ we now that

### v2

@@ 6.2 Proofs of Section˜4
$=\frac{1}{n}\sum_{i=1}^{n}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t-\frac{1}{n}\sum_{i=1}^{n}\Big(\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big)\mathds{1}_{X_{i}\notin A}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t-\frac{1}{n}\sum_{i=1}^{n}\Big(\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big)\mathds{1}_{X_{i}\notin A}.$
@@
$\leq(1+a)\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]+\frac{(1+3/a)D\log(n)^{6}}{3n}\log\Big(\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{2}}\Big)+(1+a)\tau.$
Like before, we separate $a\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$. Using Lemma˜A.1 again, we obtain
$\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big]-\int_{\mathbb{R}^{d}\setminus A}\int\int|\tilde{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y.$
@@
Collecting all other terms, inserting $\log(\sigma_{\min}^{-1})\sim\log(n)$, setting $\delta_{1}=\delta_{2}=\frac{1}{3n}$ and using the union bound leads to the result. The infimum can be used due to the presence of other nonzero terms. ∎
Preparation for the proof of Theorem˜4.3.
The proof of Theorem˜4.3 requires some additional results to bound the covering number and apply the approximation result. We further need to look carefully into the approximation result by Gühring et al. (2020).
@@
Hence we can bound $\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\leq\mathcal{N}(\frac{\tau}{\log(n)^{4}},\mathcal{M},\|\cdot\|_{\infty}).$ From Suzuki (2019, Lemma 3) we know that the covering number of the set ReLU networks of a $0,1]^{d}$, denoted by $\mathsf{NN}_{\text{ReLU}}$ cube is
@@
To consider functions on $\log(n),\log(n)]^{d}\times[0,1]$, we proceed analogously to Yakovlev and Puchkin (2025, p.46), multiplying the first $d$ coordinates of the weight matrix of the first lemma with $\log(n)$ and dividing the input vector with the same value. This has of course an impact on the bound of the weights, which will scale up by the factor $\log(n)$. Additionally, the weights of the clipping layers are of order $\log(n)^{3}$. We thus obtain
@@
The network $f_{\mathsf{NN}}$ can be chosen with at most $L\leq c\cdot\log\left(\varepsilon^{-\frac{s}{s-1}}\right)$ hidden layers and $S\leq c\cdot\varepsilon^{-\frac{d}{s-1}}\cdot\log^{2}\left(\varepsilon^{-\frac{s}{s-1}}\right)$ nonzero weights, where $c$ is a constant.
@@
As $v$ is in $C^{\infty}$, we can apply Gühring et al. (2020, Corollary 4.2) for arbitrary large $s$. However, the use of a larger $s$ will lead to a larger bound on the absolute value of the partial derivatives up to the order $s$. Since this scales up the approximation error, we need a precise quantification. The next result uses the $\log$-Sobolev inequality to bound higher derivatives of $v$.
@@
The bound on mixed derivatives follows exactly the same lines as the proofs of Lemma˜6.8 and Lemma˜6.9
Now we are ready to prove Theorem˜4.3.
Proof of Theorem˜4.3.
We scale the function down by $\sigma_{\min}^{-s-2}\log^{2}(n)$ and approximate $v\cdot\sigma_{\min}^{-s-2}\log^{2}(n)$ instead of $v$. After that, we add another layer to scale the output of the neural net up. The approximation error $\varepsilon$ will also scale up by $\sigma_{\min}^{-s-2}\log^{2}(n)$. The maximal weight in the inner network will then no longer depend on this bound, but the largest weight might increase. Hence we can set $W=C(\varepsilon^{-\frac{1}{s-2}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)).$ Inserting everything into Theorem˜4.2 and setting $\tilde{d}=d+1$, we obtain for $a<1$ with a probability of $1-\frac{1}{n}$ for $n$ big enough and $I=[-\log(n),\log(n)]^{d}$
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big]\lesssim\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t+(2+a)\tau+a(\log(n)^{4}+\log(n)^{6})$
$+\varepsilon\sigma_{\min}^{-s-2}\log^{2}(n)+\frac{a^{-1}\log(n)^{7}}{n}\varepsilon^{-\frac{\tilde{d}}{s-1}}\log\big(2\big(\tau^{-1}\log(n)^{7}L(B\vee 1)(\varepsilon^{-\frac{1}{s-1}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)\big)\big).$ (36)
Note that $B$ and $L$ will depend on $\varepsilon$, thus we are restricting feasible choices of $\varepsilon$ to choices that grow at most polynomial in $n^{-1}$. Ignoring logarithmic terms, we first solve for $a$. $\frac{a^{-1}\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\sim a$ is equivalent to $a=\Big(\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big)^{\frac{1}{2}}.$ $\tau$ can be set such that the corresponding term is of the same order, it suffices to set $\tau\sim\Big(\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big)^{\frac{1}{2}}\Big(2+\Big(\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big)^{\frac{1}{2}}\Big)^{-1}.$ This choice, as well as the choice of $\varepsilon$, influence the first term in (36) only logarithmically. Now we solve for $\varepsilon$. $\Big(\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big)^{\frac{1}{2}}\sim\varepsilon\sigma_{\min}^{-s-2}$ is equivalent to $\varepsilon\sim n^{-\frac{s-1}{\tilde{d}+2s-2}}\sigma_{\min}^{(s+2)\frac{2s-2}{\tilde{d}+2s-2}}.$ Restricting $\sigma_{\min}$ to polynomials in $n$, we ensure $\varepsilon$ is polynomial in $n$. Thus, ignoring logarithmic factors in $n$, we obtain with a probability of $1-\frac{1}{n}$ for $n$ big enough
@@
To choose $\sigma_{\min}$ such that is is optimal in the setting of Theorem˜2.1, we have to set it such that $\sigma_{\min}^{1+\alpha}\sim\Big(n^{-\frac{s-1}{d+2s-2}}\sigma_{\min}^{(s+2)\big(\frac{2s-2}{\tilde{d}+2s-2}-1\big)}\Big)^{\frac{1}{2}}$ which is equivalent to $\sigma_{\min}\sim n^{-\frac{1}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}.$ This choice is a polynomial of $n^{-1}$ and hence a feasible choice for $\sigma_{\min}$. For $n\geq 1$ and $\sigma_{\min}<1$, we know that the network approximates the Lipschitz constant of $v$ with precision $\varepsilon\sigma_{\min}^{-s-2}$. Plugging in the choice of $\sigma_{\min}$ leads to $\varepsilon\sigma_{\min}^{-s-2}\sim n^{-\frac{2+2\alpha}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}\leq 1.$ Hence we obtain for the Lipschitz constant $\hat{L}_{t}$ of $\hat{v}_{t}$ $e^{\int_{0}^{1}\hat{L}_{t}\;\mathrm{d}t}\leq e^{\int_{0}^{1}\Gamma_{t}+1\;\mathrm{d}t}\leq e^{C+1}.$ Note that we cannot let $s\rightarrow\infty$, as this will blow up the constant. However, for every fixed $\eta>0$, we can choose $s$ such that $s=\big\lceil(5+2\alpha)\frac{\tilde{d}}{\eta}+1\big\rceil.$ This leads to the final choice $\sigma_{\min}\sim n^{-\frac{1}{\tilde{d}+4\alpha+4+\eta}}$ Now we use the results of Lemma˜6.5, set $\delta=\frac{1}{2n}$ and use the union bound. Combining this with Theorem˜2.1, Theorem˜3.5, (Kunkel and Trabs, 2025b, Theorem 3.3) and $\tilde{d}=d+1$, we obtain with a probability of $1-\frac{1}{n}$
@@
Appendix A Additional proofs of Section˜6 and helper results
Lemma A.1.
Proof of Lemma˜A.1.
Setting $\tilde{v}_{t}=v_{t}$, we obtain
$0=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big]+C\quad\Longleftrightarrow\quad C=-\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big].\qed$
Proof of Lemma˜6.1.
@@ Appendix A Additional proofs of Section˜6 and helper results
$\frac{\partial}{\partial x_{i}}v^{j}_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ji}.\qed$
Proof of Lemma˜6.2.
@@
$\int_{0}^{t^{*}}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\Big(\gamma t^{\gamma-1}$ $-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\Big|\;\mathrm{d}t\int_{0}^{t^{*}}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|\;\mathrm{d}t+\int_{0}^{t^{*}}\Big|\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\Big|\;\mathrm{d}t.$
@@
For the second integral, we can bound the Covariance term by $\big|\operatorname{Cov}(Y^{\cdot,t})_{ij}\big|\leq C.$ Now we get
@@
$\leq C\int_{0}^{t^{*}}\Big|\gamma t^{\gamma-1}\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big|\;\mathrm{d}t+C\int_{0}^{t^{*}}\Big|\frac{\sigma_{t}^{\prime}t^{2\gamma}}{\sigma^{3}_{t}}\Big|\;\mathrm{d}t\leq\vartheta^{2}C\int_{0}^{t^{*}}\gamma t^{\gamma-1}\;\mathrm{d}t+\vartheta^{2}C\int_{0}^{t^{*}}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta)).\qed$
Proof of Lemma˜6.5.
@@
$\int_{\mathbb{R}^{d}\setminus A}p^{*}(y)\;\mathrm{d}y\leq\frac{\int_{\mathbb{R}^{d}\setminus A}|y|^{2}\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}\leq e^{2L}\Big(2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big)^{d}.$
@@
$0\leq\int_{\log(n)}^{\infty}y_{1}^{2}\exp\Big(-\frac{y_{1}^{2}}{2}\Big)\;\mathrm{d}y=\Big[-y_{1}\exp\Big(-\frac{y_{1}^{2}}{2}\Big)\Big]_{\log(n)}^{\infty}-\int_{\log(n)}^{\infty}\exp\Big(-\frac{y_{1}^{2}}{2}\Big)\;\mathrm{d}y_{1}\leq\log(n)\exp\Big(-\frac{\log(n)^{2}}{2}\Big),$
@@
$e^{2L}\Big(2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big)^{d}\lesssim\Big(\log(n)\exp\Big(-\frac{\log(n)^{2}}{2}\Big)\Big)^{d}.$
@@
$\mathbb{E}_{X_{i}}[|B|]\lesssim\log(n)^{6}\Big(\log(n)\exp\Big(-\frac{\log(n)^{2}}{2}\Big)\Big)^{d}.$
Now Markovs inequality yields
$\mathbb{P}(B\geq n^{-\frac{1}{2}})\lesssim\log(n)^{6}\Big(\log(n)\exp\Big(-\frac{\log(n)^{2}}{2}\Big)\Big)^{d}n^{\frac{1}{2}},\quad\mathbb{P}(B\leq-n^{-\frac{1}{2}})\lesssim\log(n)^{6}\exp\Big(-\frac{\log(n)^{2}}{2}\Big)\Big)^{d}n^{\frac{1}{2}}.$
@@
$\log(n)^{6}\Big(\log(n)\exp\Big(-\frac{\log(n)^{2}}{2}\Big)\Big)^{d}n^{\frac{1}{2}}\lesssim\frac{1}{3n}\quad\Longleftarrow\quad n\gtrsim e^{3(d+7+\frac{1}{2})},$
@@
Proof of Lemma˜6.6.
From the proof of Theorem˜4.2 and as $\log(\sigma_{\min}^{-1})\sim\log(n)$ we now that